Gionino's Pizzeria is considering opening a drive-through window for customer pickups. It is estimated that at peak times (5-9pm), customers will arrive at a rate of 10.3 per hour. The clerk who will staff the window can serve customers at the rate of 1 every 4 minutes.

For the peak times, what is the average waiting time in line (in minutes)?

Give your answer to 2 decimal places and do NOT list minutes. For example, answer like 3.24
Answer #1

Let's define the variables:

  • a = Arrival Rate
  • b = Service Rate

a = 10.3/hour

b = 1/4 per min= 15/hour


Average customers waiting in Line = (10.3)^2/(15(15-10.3)) = 1.068 customers
Average waiting time in line for customers = 1.068 / 10.3 = 0.10368 hours = 6.22